OpenStudy (anonymous):

P = 2L + 2W solve for L

OpenStudy (anonymous):

subtract 2W from both sides P - 2W = 2L + 2W - 2W P - 2W = 2L finally divide both sides by 2 L = (P - 2W) / 2